You are here: Foswiki>Commissioning Web>Logs>13Sep2012_Log (14 Sep 2012, JuanGuerra)Edit Attach

UT 20120913 AO#2 Commissioning

Observer: R. Briguglio, A. Puglisi, E. Pinna (INAF)
Telescope Operator: D. Gonzalez(LBTO)
Additional AO Support: JCGuerra (LBTO)
Telescope Support: (LBTO)
Instrument Support: J.Morris (LBTO)

Summary

Startup of wfssx.failed. The BCU proceses was hung. This is known problem. The systems was restarted doing stop/start (w_stop/w_start) ASM_SX set ok.

Analysis of the effect of the tree-house ventilation on the images. The images are taken with fast frames with the IRTC at different setup (tree-house ventillation on/off, EL breaks(set/rel) with HBS ON .

Acquisition and analysis of slope null for LUCIFER.

The shell RIP during a close loop, the reason it is not clear.

Problem with the pic controller. This failure stop the measurements for about 2 hours. This problem has to be checked.

Results will be in the coming Arcetri team progress report.

Details

4D.

Nothing to report. Interferometer working properly. It is still used to acquire interferograms images.

Secondary

The secondary is working at open and close loop and it behaves is good, the software is working ok, and we had RIP shell dueing the close loop, the reason of the RIP is unclear. time 17:10:40

Left W# unit

Acquisition of data to analyse the effect of the tree-house ventilation on performances. The data will be reduced and the results include in the next Arcetri progress report.

Acquisition of the slope-null to determine how much the wfs can reach.

Problem with the pic controller that stop the measurements for about 2 hours.

-- JuanGuerra - 13 Sep 2012
Topic revision: r2 - 14 Sep 2012, JuanGuerra
This site is powered by FoswikiCopyright © by the contributing authors. All material on this collaboration platform is the property of the contributing authors.
Ideas, requests, problems regarding Foswiki? Send feedback